# RAGentA: Multi-Agent Retrieval-Augmented Generation for Attributed Question Answering

### v1

@@ Abstract.
We present RAGentA, a multi-agent retrieval-augmented generation (RAG) framework for attributed question answering (QA). With the goal of trustworthy answer generation, RAGentA focuses on optimizing answer correctness, defined by coverage and relevance to the question and faithfulness, which measures the extent to which answers are grounded in retrieved documents. RAGentA uses a multi-agent architecture that iteratively filters retrieved documents, generates attributed answers with in-line citations, and verifies completeness through dynamic refinement. Central to the framework is a hybrid retrieval strategy that combines sparse and dense methods, improving Recall@20 by 12.5% compared to the best single retrieval model, resulting in more correct and well-supported answers. Evaluated on a synthetic QA dataset derived from the FineWeb index, RAGentA outperforms standard RAG baselines, achieving gains of 1.09% in correctness and 10.72% in faithfulness. These results demonstrate the effectiveness of the multi-agent architecture and hybrid retrieval in advancing trustworthy QA.
@@ 1. Introduction
Large Language Models (LLMs) are increasingly applied to tasks such as multi-step reasoning (Wei et al., 2022), summarization (Zhang et al., 2024), and open-domain QA (Singhal et al., 2025). Nevertheless, LLMs often generate hallucinations, meaning factually incorrect or misleading outputs (Huang et al., 2025). Because users currently have limited means to verify LLM-generated answers, hallucinated information becomes especially problematic. This lack of verifiability poses a significant challenge for deploying LLMs in high-stakes domains such as science, healthcare, and law, where factuality and trust are essential.
RAG addresses this issue by combining traditional information retrieval with LLMs (Lewis et al., 2020). Relevant documents are retrieved from external sources and used to guide the model’s generation. This approach reduces the likelihood of unsupported claims by grounding answers in curated, up-to-date information. Building on this foundation, Gao et al. (2023b) show that RAG can also enable attribution, for example via in-line citations. This allows users to verify the origins of generated content and increases trustworthiness.
@@
We propose RAGentA, a collaborative multi-agent RAG framework that improves answer faithfulness through fine-grained attribution, especially for multi-source questions.
@@ 3. Synthetic Dataset Generation
We constructed a diverse synthetic QA benchmark dataset for the evaluation of our RAGentA framework using the DataMorgana platform (Filice et al., 2025). The dataset comprises 500 QA pairs, each associated with multiple supporting evidence paragraphs. Detailed statistics of the dataset, including the distribution across various question and user categories, are presented in Table 1.
@@ 4.1. Hybrid Retrieval System
The first stage of our framework is a hybrid retrieval system that combines both dense and sparse retrieval to ensure broad and contextually relevant document coverage. Sparse retrieval, based on BM25, is a widely adopted model that relies on exact term matching (Robertson and Walker, 1994; Robertson et al., 1994). Our semantic retrieval is powered by dense embeddings using the intfloat/e5-base-v2 model (E5) with Pinecone as the vector store to capture nuanced semantic relationships (Wang et al., 2024).
@@ 4.2.1. Agent-1: Predictor
Following document retrieval, Agent-1 generates query-specific answers for each retrieved document, forming Document-Query-Answer triplets. This agent employs an LLM to analyze each document in the context of the given query and generates an answer based on its content.
@@ Acknowledgements.
- Lewis et al. (2020) Patrick Lewis, Ethan Perez, Aleksandra Piktus, Fabio Petroni, Vladimir Karpukhin, Naman Goyal, Heinrich Küttler, Mike Lewis, Wen-tau Yih, Tim Rocktäschel, Sebastian Riedel, and Douwe Kiela. 2020. Retrieval-Augmented Generation for Knowledge-Intensive NLP Tasks. In Advances in Neural Information Processing Systems, Vol. 33. Curran Associates, Inc., Red Hook, NY, USA, 9459–9474. https://doi.org/10.5555/3495724.3496517
@@
- Robertson and Walker (1994) S. E. Robertson and S. Walker. 1994. Some simple effective approximations to the 2-Poisson model for probabilistic weighted retrieval. In ACM SIGIR’94 (Dublin, Ireland). Springer-Verlag, Berlin, Heidelberg, 232–241.
@@
- Wei et al. (2022) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, Brian Ichter, Fei Xia, Ed H. Chi, Quoc V. Le, and Denny Zhou. 2022. Chain-of-thought prompting elicits reasoning in large language models. In Proceedings of the 36th International Conference on Neural Information Processing Systems (New Orleans, LA, USA) (NIPS ’22). Curran Associates Inc., Red Hook, NY, USA, Article 1800, 14 pages.

### v2

@@ Abstract.
We present RAGentA, a multi-agent retrieval-augmented generation (RAG) framework for attributed question answering (QA) with large language models (LLMs). With the goal of trustworthy answer generation, RAGentA focuses on optimizing answer correctness, defined by coverage and relevance to the question and faithfulness, which measures the extent to which answers are grounded in retrieved documents. RAGentA uses a multi-agent architecture that iteratively filters retrieved documents, generates attributed answers with in-line citations, and verifies completeness through dynamic refinement. Central to the framework is a hybrid retrieval strategy that combines sparse and dense methods, improving Recall@20 by 12.5% compared to the best single retrieval model, resulting in more correct and well-supported answers. Evaluated on a synthetic QA dataset derived from the FineWeb index, RAGentA outperforms standard RAG baselines, achieving gains of 1.09% in correctness and 10.72% in faithfulness. These results demonstrate the effectiveness of our multi-agent RAG architecture and hybrid retrieval strategy in advancing trustworthy QA with LLMs.
@@ 1. Introduction
LLMs are increasingly applied to tasks such as multi-step reasoning (Wei et al., 2022), summarization (Zhang et al., 2024), and open-domain QA (Singhal et al., 2025). Nevertheless, LLMs often generate hallucinations, meaning factually incorrect or misleading outputs (Huang et al., 2025). Because users currently have limited means to verify LLM-generated answers, hallucinated information becomes especially problematic. This lack of verifiability poses a significant challenge for deploying LLMs in high-stakes domains such as science, healthcare, and law, where factuality and trust are essential.
RAG addresses this issue by combining traditional information retrieval with LLMs (Lewis et al., 2020). Relevant documents are retrieved from external sources and used to guide the model’s generation. This approach reduces the likelihood of unsupported claims by grounding answers in curated, up-to-date information. Building on this foundation, Gao et al. (2023b) show that RAG can also enable attribution, for example via in-line citations. This allows users to verify the sources of generated content and increases trustworthiness.
@@
We propose RAGentA, a collaborative multi-agent RAG framework that improves answer faithfulness through fine-grained citations, especially for multi-source questions.
@@ 3. Synthetic Dataset Generation
We constructed a diverse synthetic QA benchmark for the evaluation of our RAGentA framework using the DataMorgana platform (Filice et al., 2025). The dataset comprises 500 QA pairs, each associated with multiple supporting evidence paragraphs. Detailed statistics of the dataset, including the distribution across various question and user categories, are presented in Table 1.
@@ 4.1. Hybrid Retrieval System
The first stage of our framework is a hybrid retrieval system that combines both dense and sparse models to ensure broad and contextually relevant document coverage. Sparse retrieval, based on BM25, is a widely adopted model that relies on exact term matching (Robertson and Walker, 1994; Robertson et al., 1994). Our semantic retrieval is powered by dense embeddings using the intfloat/e5-base-v2 model (E5) with Pinecone as the vector store to capture nuanced semantic relationships (Wang et al., 2024).
@@ 4.2.1. Agent-1: Predictor
Following document retrieval, Agent-1 generates query-specific answers for each retrieved document, forming Query-Document-Answer triplets. This agent employs an LLM to analyze each document in the context of the given query and generates an answer based on its content.
@@ Acknowledgements.
- Lewis et al. (2020) Patrick Lewis, Ethan Perez, Aleksandra Piktus, Fabio Petroni, Vladimir Karpukhin, Naman Goyal, Heinrich Küttler, Mike Lewis, Wen-tau Yih, Tim Rocktäschel, Sebastian Riedel, and Douwe Kiela. 2020. Retrieval-Augmented Generation for Knowledge-Intensive NLP Tasks. In Advances in Neural Information Processing Systems, Vol. 33. Curran Associates, Inc., Red Hook, NY, USA, 9459–9474. https://dl.acm.org/doi/abs/10.5555/3495724.3496517
@@
- Robertson and Walker (1994) S. E. Robertson and S. Walker. 1994. Some simple effective approximations to the 2-Poisson model for probabilistic weighted retrieval. In ACM SIGIR’94 (Dublin, Ireland). Springer-Verlag, Berlin, Heidelberg, 232–241. https://doi.org/10.1007/978-1-4471-2099-5_24
@@
- Wei et al. (2022) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, Brian Ichter, Fei Xia, Ed H. Chi, Quoc V. Le, and Denny Zhou. 2022. Chain-of-thought prompting elicits reasoning in large language models. In Proceedings of the 36th International Conference on Neural Information Processing Systems (New Orleans, LA, USA) (NIPS ’22). Curran Associates Inc., Red Hook, NY, USA, Article 1800, 14 pages. https://dl.acm.org/doi/10.5555/3600270.3602070